OpenStudy (anonymous):

We are given the polar equation r = cos(4θ). If 0 < θ < π and r = 0, then θ =?

OpenStudy (anonymous):

if r is 0 then there is no radius if there's no radius there's no circle. I think something's missing from the problem.

OpenStudy (anonymous):

the graph is a rose and is symmetrical on x and y axis. So when r=0 theta would not be there?

OpenStudy (anonymous):

cos(90)=0 so,4θ=90 θ=22.5= π /8

OpenStudy (anonymous):

there is more than one theta, there are four

OpenStudy (anonymous):

So basically we have\[0=\cos(4\theta)\] which means \[4\theta=(n+1/2)*\pi\]where n is an integer. However \[0<\theta<\pi\] which means that\[\theta=\pi/8\]

OpenStudy (anonymous):

so I am right

OpenStudy (anonymous):

Actually I lied there are more solutions. Just a sec.

OpenStudy (anonymous):

where did you get 4theta=(n+1/2)*pi from?

OpenStudy (anonymous):

you can multiply by n where n=1,2,3...8

OpenStudy (anonymous):

\[\Theta=(n+1/2)*\pi/4\] So the solutions are:\[\Theta=\pi/8, 3\pi/8, 5\pi/8, 7\pi/8\]

OpenStudy (anonymous):

I got that from the fact that cos((n+1/2)*pi)=0. Just like sin(npi)=0.

OpenStudy (anonymous):

cos(pi/2)=cos(3pi/2)=cos(5pi/2)=...=0

OpenStudy (anonymous):

ok, that makes sense but why would you set 4theta to 0 instead of cos(4theta)?

OpenStudy (anonymous):

I didn't. I set 4theta to the value which makes cos(4theta)=0.

OpenStudy (anonymous):

I understand that but why?

OpenStudy (anonymous):

cos(90)=0 so,4θ=90 θ=22.5= π /8

OpenStudy (anonymous):

Because cos(4theta)=r=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Does that all make sense? Trig functions get tricky because you have to be very careful that you operate within the given domain and don't miss any solutions.

OpenStudy (anonymous):

kind of, I understand trig part but the reason why the solution is taken like that is confusing.

OpenStudy (anonymous):

Can I explain something better? Which part specifically is confusing?

OpenStudy (anonymous):

the answer θ=22.5= π /8

OpenStudy (anonymous):

why cant you use tantheta=y/x and solve for x and y based on the r given. so, x=rcostheta and y=rsintheta. then solve for theta?

OpenStudy (anonymous):

Well you are given r=0 and you aren't given information about cartesian coordinates so I would recommend doing it in polar coordinates.

OpenStudy (anonymous):

Think of it like solving the equation y=sinx and they tell you y=0. What do you do? You solve it similarly: x=npi. And then you would worry about the domain for x.

OpenStudy (anonymous):

ok I think I understand, thanks

OpenStudy (anonymous):

Ok good luck with your studies. ;p

OpenStudy (anonymous):

sorry i answered that quite tired, basically what i meant was that if r=0 theta would equal 0 so you need to find the answers of theta that would allow the cos function go to 0. 0=Cos (theta) values theta pi/2 and 3pi/2 but since there is no radius it is simply the axis and you're at the origin, but you still need those values in order the trace the rose curve. Hopefully that helps a bit more. =)